# Tailoring anisotropy in kirigami metamaterial skins with pop-up folding hinges

### v1

@@ 4.1 Kinematic analysis
To track the evolution of the pop-ups with $\theta$, we derive formulas for their heights. To do so, we introduce some auxiliary parameters $a_{i}$ and $b_{i}$ for each pop-up, which define right triangles in which the third dimension is $\ell_{i}$ and one of the angles is $\lambda_{i}$, as shown in Fig. 5A and as indicated in the following formulas:
@@
To calculate the pop-up heights, we consider a deformed configuration of the unit cell (either Fig. 5B or C, where quantities of interest are only shown for the pop-up of the 1 type). We then concentrate on the triangle formed by the pop-up height $h_{i}$, the segment of length $b_{i}$, which is now also coming out of plane, and its projection on the $x$-$y$ plane. Using Pythagoras’ theorem on this triangle, we determine that the height of a generic pop-up can be derived from
@@
Provided that all panels within the unit rotate of the same amount, which is a requirement to have the desired global mechanism, this equation can be applied to determine the height of any pop-up, independently of where it is located within the unit. In the following, pop-up heights are normalized as follows: ¯$\bar{\ell}_{1}=\ell_{1}/L_{y}$, ¯$\bar{\ell}_{2}=\ell_{2}/L_{x}$, ¯$\bar{\ell}_{3}=\ell_{3}/L_{y}$, ¯$\bar{\ell}_{4}=\ell_{4}/L_{x}$. The evolution of the pop-up heights with $\theta$ for the example geometry is shown in the middle panel of Fig. 5D. In general, these curves always increase; while they start steep, they tend to flatten as $\theta$ reaches its limit.
$A_{i}=h_{i}a_{i}\sin{(\lambda_{i}-\theta)}.$ (7)
This quantity, albeit not the most accurate representation of pop-up area, allows us to monitor the relative opening between pop-ups — something that is of key importance, even more than the relative pop-up height, to obtain anisotropic friction.
Finally, we use these formulas to perform kinematic simulations, that allow us to visualize the evolution of the unit cell with $\theta$. Figs. 5A–C represent three instances of one such simulation. To obtain these snapshots, we apply planar rotations to the octagonal panels; the pop-ups, on the other hand, are obtained by using the pop-up height equation (Eq. 6), and by connecting the pop-up peaks to the proper points on the octagons.
4.2 Further restriction of the parameter space
Even if we restrict ourselves to patterns with $\ell_{1}=\ell_{3}$ and $\ell_{2}=\ell_{4}$, not all combinations of the remaining parameters are valid. In fact, some combinations yield unwanted overlaps between panels and hinges. To avoid overlaps, we set limits on these parameters, as indicated in the following; these limits are derived from geometrical considerations on the unit cell in Fig. 5A. Slit parameters $w_{x}$ and $w_{y}$ are chosen such that
@@ 4.2 Further restriction of the parameter space
Fig. 6: Kinematics-guided design space exploration, part 1 (part 2 is in Fig. 7). (A) Parametric study in which we set fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$, fold angles $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$, and let slit openings $\bar{w}_{x}$ and $\bar{w}_{y}$ vary. (B) Parametric study in which we set slit openings $\bar{w}_{x}=\bar{w}_{y}=0$, fold angles $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$, and let fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{3}$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}$ vary. Throughout the figure, ($\square$-1) subfigures indicate the max horizontal strain $\varepsilon_{x}$ for each parameter combination, which we consider as reference strain for the other maps; ($\square$-2) subfigures show the tangential strain $\varepsilon_{y}$, ($\square$-3) the effective Poisson’s ratio, while subfigures ($\square$-4)–($\square$-7) show the normalized pop-up heights and pop-up opening areas; if several pop-up types behave identically, subfigures are grouped as indicated in their individual titles. Blackened-out regions of the maps correspond to parameter combinations that yield panel overlaps and, therefore, unphysical patterns. In the Poisson’s ratio map, gray regions are such that the Poisson’s ratio itself cannot be defined since $\varepsilon_{x}=0$. (A-6), (A-7) Kinematic simulations of sample geometries with parameters corresponding to the (A) study, and with $\bar{w}_{x}=\bar{w}_{y}=0$ and $\bar{w}_{x}=0.3$, $\bar{w}_{y}=0$, respectively. (B-8) Kinematic simulation of sample geometry with parameters corresponding to the (B) study, and with ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.2$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.7$.
Fig. 7: Kinematics-guided design space exploration, part 2 (subfigure labels continue from Fig. 6). (C) Parametric study in which we set slit openings $\bar{w}_{x}=\bar{w}_{y}=0$, fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$, and let fold angle $\lambda_{2}$ vary independently from $\lambda_{1}=\lambda_{3}=\lambda_{4}$. (D) Parametric study in which we set slit openings $\bar{w}_{x}=\bar{w}_{y}=0$, fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.2$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.7$, and let fold angle $\lambda_{2}$ vary independently from $\lambda_{1}=\lambda_{3}=\lambda_{4}$. Subfigure labeling follows the same convention as in Fig. 6. (C-8) Kinematic simulation of sample geometry with parameters corresponding to the (C) study, and with $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.39$ and $\lambda_{2}=1.15$. (D-10) Kinematic simulation of sample geometry with parameters corresponding to the (D) study, and with $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.45$, $\lambda_{2}=0.85$.
4.3 Kinematics-guided parametric study
In study A (Fig. 6A), we probe the effects of the slit openings $\bar{w}_{x}$ and $\bar{w}_{y}$, while fixing hinge lengths to ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$ and angles to $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$. In Fig. 6A-1, we notice that a large portion of the design space, the black region, is not accessible. In the accessible region, we can see that patterns with $\bar{w}_{y}>0.17$ do not produce any positive strain and only contract along $x$ for all admissible values of $\theta$; these designs are of no interest to us here. We can also see that $\bar{w}_{x}$ does not affect the maximum strain along $x$. From Figs. 6A-2,A-3, where we plot the $\varepsilon_{y}$ strain and Poisson’s ratio at the $\theta$ value of maximum positive $\varepsilon_{x}$, we can see that the slit opening parameters allow to vastly tune the Poisson’s ratio of the pattern. In particular, low and high $\bar{w}_{x}$ values yield negative and positive Poisson’s ratios, respectively. For this choice of parameters, the folding hinges are all identical. Thus, Figs. 6A-4,A-5 are representative of all pop-up heights and opening areas, respectively. We can see that the pop-up height correlates with $\varepsilon_{x}$, while the opening area has a maximum for values of $\bar{w}_{y}$ around 0.1. To conclude this first study, we select two representative unit cells corresponding to combinations of parameters that yield symmetric designs with negative and positive Poisson’s ratios. Their kinematic evolution, for $\theta$ values until the maximum $\varepsilon_{x}$ is reached, is shown in Fig. 6A-6,A-7, where we also detail the exact parameters needed to obtain both patterns.
In study B (Fig. 6B), we instead investigate the effects of the hinge lengths. Owing to the previously-determined constraint on hinge lengths, we select as independent parameters ¯$\bar{\ell}_{1}=\bar{\ell}_{3}$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}$, while setting slit openings $\bar{w}_{x}=\bar{w}_{y}=0$ and fold angles $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$. Figs. 6B-1–B-3 show similar trends to the corresponding figures of the A study, highlighting that hinge lengths can also be used to tune the Poisson’s ratio — since hinge lengths, just like slit opening parameters, allow to change the location of the rotation points of the panels. More interesting here are the maps for pop-up heights and opening areas in Figs. 6B-4–B-7. These maps show that the height of pop-ups 2 and 4 can be maximized by selecting large lengths for these hinges and small lengths for the others. On the other hand, to maximize the opening areas of pop-ups 2 and 4, one needs to choose ¯$\bar{\ell}_{1}=\bar{\ell}_{3}\approx 0.3$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}\approx 0.7$ — something that would have been hard to guess without a systematic parametric study. The kinematic evolution of a non-auxetic unit with ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.2$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.7$ is shown in Fig. 6B-8, where we can appreciate how large the pop-ups of the 2 and 4 types can be. Conversely, these maps show that pop-ups cannot be as pronounced in units with negative $\nu$.
In studies C and D (Figs. 7C,D), we instead concentrate on the effects of the hinge angles, with an eye towards obtaining patterns with drastically-asymmetric pop-ups. We therefore choose as independent parameters hinge angles $\lambda_{2}$ and $\lambda_{1}=\lambda_{2}=\lambda_{3}$ — something that would cause the resulting patterns to have frictional asymmetry along the $x$ axis. In study C, as shown in Figs. 7C-1–C-3, we concentrate on an asymmetric pattern with negative Poisson’s ratio. The effects of hinge angles on the extensibility of the cell are minimal, as long as these angles are not so small to cause a premature, complete folding of the hinges. The colormaps for pop-up heights and opening areas in Figs. 7C-4–C-7 show that the region that yields more asymmetry in favor of the pop-up of the 2 type corresponds to large $\lambda_{2}$ and to small values of all other hinge angles. Thus, we choose the geometry shown in Fig. 7C-8, with $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.39$ and $\lambda_{2}=1.15$, as the representative asymmetric one with negative Poisson’s ratio; we did not choose a smaller value for $\lambda_{1}$ since this could have limited the overall extensibility of the unit. As illustrated by this kinematic simulation, the pop-up of the 2 type is much larger and much more open than the others at the maximum strain level — which is exactly what we are looking for. In study D, we instead concentrate on obtaining similar effects in geometries with positive Poisson’s ratio that are morphologically similar to the one shown in Fig. 6B-8, i.e., with slit openings $\bar{w}_{x}=\bar{w}_{y}=0$ and fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.2$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.7$. From the maps in Figs. 7D-4–D-9, we can see that the maximum asymmetry in pop-up opening areas is achieved with $\lambda_{2}=0.85$ and all other angles $\approx 0.2$. Yet, Figs. 7D-1–D-3 show that there is a tradeoff between asymmetry and maximum cell extension. Thus, as representative asymmetric geometry with positive Poisson’s ratio, we select one with $\lambda_{1}=\lambda_{3}=\lambda_{4}=0.45$, $\lambda_{2}=0.85$, and show its kinematic evolution in Fig. 7D-10. This simulation allows us to appreciate that designs with positive $\nu$ can produce much more pop-up asymmetry with respect to those with negative $\nu$.
5 Response of representative designs
Leveraging our previous design space explorations, we now analyze the response of a few representative symmetric and asymmetric geometries. First and foremost, we validate our analytical kinematics results for patterns that display mechanism-like behavior, via experiments and numerical simulations. We also simulate the behavior of a geometry that, according to our previous analysis, does not feature a mechanism. Then, we use numerical simulations alone to extract some preliminary information on the force-deformation response of these patterns. The selected representative patterns are listed in Table 1, where we indicate the figure(s) they appear in and their design parameters.
$\bar{w}_{x}$ $\bar{w}_{y}$$\bar{\ell}_{1}$$\bar{\ell}_{2}$$\bar{\ell}_{3}$$\bar{\ell}_{4}$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$
Fig. 9A 0 0 0.3 0.3 0.3 0.3 $\pi/4$ $\pi/4$ $\pi/4$ $\pi/4$
Fig. 6A-6
Fig. 9B 0.3 0 0.3 0.3 0.3 0.3 $\pi/4$ $\pi/4$ $\pi/4$ $\pi/4$
Fig. 6A-7
Fig. 9C 0 0 0.3 0.3 0.3 0.3 0.39 1.15 0.39 0.39
Fig. 7C-8
Fig. 9D 0 0 0.2 0.7 0.2 0.7 0.45 0.85 0.45 0.45
Fig. 7D-10
Fig. 9E 0 0 0.2 0.4 0.2 0.2 $\pi/6$ $\pi/4$ $\pi/6$ $\pi/6$
Table 1: Geometrical parameters of the representative designs we select for further analysis, indicating the figures where such patterns appear.
5.1 Fabrication and experiments
Representative unit cells are fabricated out of 0.25 mm-thick PETG sheets, via laser cutting (by setting our 80 W Epilog Fusion Pro 32 machine to 45% speed, 50% power and 100% frequency). As shown in Fig. 8A, which illustrates one such unit cell, fold lines are laser-engraved (500 DPI, 100% speed, 50% power). To avoid tearing of the sheet at the meeting points of fold lines, we add stress-relieving circular cuts as in Ref. [37].
Fig. 8: Fabricated sample and experimental setup. (A) Top view of the flat, undeformed configuration of the specimen of one of our representative patterns. (B) Specimen in (A), subjected to a $\theta=20^{\text{o}}$ actuation; the specimen is kept in this configuration with pieces of double-sided tape under each panel. We use a caliper to measure (C) the extension of the unit and (D) the height of the pop-ups for different actuation angles.
@@ 5.1 Fabrication and experiments
For all unit cells, we set $L_{x}=L_{y}=4.5\,$cm. Our measurements on these specimens are only intended to characterize how their shape changes during their mechanism motion, and therefore only probe the kinematics of the unit cell. To perform these measurements, we manually actuate the specimen to a desired opening angle with the aid of laser cut wedges, as shown in Fig. 8B; if, as in the example, slit angles are nonzero to begin with, wedges account for the sum of the slit angle and the desired actuation angle. The octagonal panels are then taped to a flat surface with double-sided tape. As shown in Figs. 8C,D, we use a caliper to measure the overall cell extension (and, therefore, in-plane strains) and pop-up heights for various $\theta$ values. To account for specimen positioning and measurement inaccuracies, we perform each measurement three times, re-positioning the specimen each time. As a note, some plastic deformation and occasional tearing at the hinges is inevitable in these specimens, owing to the relatively small thickness difference between panels and fold lines. To weaken the fold lines and allow the specimens to better follow the desired mechanism motion, we precondition specimens by repeatedly folding panels about all fold lines.
5.2 Finite element model

### v2

@@ B.1 Steps for the pop-up height formulas
Here, we report some intermediate steps needed to obtain formulas for the pop-up heights $h_{i}$ as a function of $\theta$. We introduce some auxiliary parameters $a_{i}$ and $b_{i}$ for each pop-up, which define right triangles in which the third dimension is $\ell_{i}$ and one of the angles is $\lambda_{i}$, as shown in Fig. 3(a) and as indicated in the following formulas:
@@
To calculate the pop-up heights, we consider a deformed configuration of the unit cell (either Fig. 3(b) or (c), where quantities of interest are only shown for the pop-up of the 1 type). We then concentrate on the triangle formed by the pop-up height $h_{i}$, the segment of length $b_{i}$, which is now also coming out of plane, and its projection on the $x$-$y$ plane. Using Pythagoras’ theorem on this triangle, we determine that the height of a generic pop-up can be derived from
@@
B.2 Further restriction of the parameter space
Some combinations of parameters yield unwanted overlaps between panels and hinges. To avoid overlaps, we set limits on these parameters, derived from geometrical considerations on the unit cell in Fig. 3(a). Slit parameters $w_{x}$ and $w_{y}$ are chosen such that
@@ B.2 Further restriction of the parameter space
Appendix C Fabrication details
Representative unit cells and tubular structures are fabricated via laser cutting and manual assembly. For all samples, we set the unit cell panel lengths to $L_{x}=L_{y}=2.5\,$cm. All samples feature a thin-sheet layer with fold lines and cuts, laser cut out of 0.05 mm-thick PETG sheets. The cut pattern of one of our unit cells is shown in Fig. 10(a); to facilitate folding along desired fold lines, we weaken some regions by dotted cuts, with dots approximately 1 mm distant from each other. These dotted cuts also act as stress relieving locations to prevent significant tearing at the fold lines (while some tearing after heavy usage is inevitable).
Fig. 10: Fabrication of unit cells and tubular structures. (a) Laser cut pattern for the thin substrate of a single unit cell of the AN type. (b) Laser cut pattern for the stiffening panels of the same unit cell; we directly laser cut panels with double-sided tape attached to them for easier assembly. (c) Assembled unit cell. (d) Laser cut pattern for the thin substrate of a tubular structure with AN geometry. The two vertical strips without cuts at the left and right edges are the locations where the octagonal end caps are bonded. The tabs on the bottom edge, where more double sided tape is located, are used to close the tube once the skin is wrapped around the end caps. (e) One of our fabricated tubes, with a drawing of one of the end caps.
All of our unit cells and some of our tubular structures also feature stiffening panels, that allow the structures to closely follow the behavior expected from our kinematic calculations. To create these panels and to facilitate the process of bonding them to the thin substrate, we apply a double sided tape sheet (3M 300LSE) onto a 0.5 mm-thick PETG sheet, and laser cut them together according to patterns as the one in Fig. 10(b). We then peel the other protective layer of the double sided tape and manually position the panels onto the substrate, obtaining unit cells as the one in Fig. 10(c).
Fabricating tubular structures requires some additional steps. Due to the presence of stiffening panels, our skins are only suitable to create tubes with polygonal, rather than circular, cross sections. Here, we choose to create tubes with octagonal sections, and this can be achieved by considering four cells along the circumferential direction. We instead choose five cells along the axial direction. We fabricate tubes of the four representative geometries in Table 1; for each geometry, we create a tube with stiffening panels and one with only the thin substrate, to compare their mechanics and the effects of these panels.
To facilitate bonding of octagonal end caps to our tube, we add strips without cuts to our laser cut pattern, as shown in Fig. 10(d). To be able to close the tubes once they are wrapped around the end caps, we add tabs to one of the two skin edges running along the axis of the tube, as shown in Fig. 10(d).
The tubes are assembled in two steps. First, we use double sided tape to bond the end caps to the skin. The end caps inevitably alter the way boundary unit cells behave, by constraining their deformation. The end caps have rectangular openings needed to fix them to our UTS via custom grippers, as discussed in D; it is important that the rectangular holes on the two end caps of the same tube are aligned with each other. The caps also feature an engraved rectangular area on their internal face, to easily position the custom gripper. After bonding the skin to the end caps, we use double sided tape to bond the tabs to the hinges on the opposite side of the skin, as to close the tubes without altering the kinematics of the unit cells at the bonded edge. A tube resulting from our fabrication process is shown in Fig. 10(e).
Appendix D Experimental setups
In this appendix, we report details on our experimental procedures and the setups used throughout this work.
Fig. 11: Experimental setups. (a) Image of a unit cell, taped to a surface and opened at a desired angle $\theta=20^{\text{o}}$ via a wedge. Extension and pop-up heights are measured via a caliper. (b) Experimental setup to evaluate the morphing mechanics of tubes. (c) Drawing of the custom grippers. (d) Photos showing how one of the grippers is anchored to the tube via insertion, rotation and pulling. (e) Experimental setup to evaluate the anisotropic friction of a tube against a rough surface; $d$ is the distance between the axis of the tube and the surface. (f) Drawing of the actuation bar used to keep the tube at a desired displacement (2 cm in this case) and to attach the open tube to the UTS. (g) Photos showing how the actuation bar is used to elongate the tube at a desired displacement.
D.1 Experiments to validate unit cell kinematics
@@
Our measurements on single unit cells are only intended to characterize how their shape changes during their mechanism motion, and therefore only probe their kinematics. To perform these measurements, we manually actuate the specimen to a desired opening angle with the aid of laser cut wedges, as shown in Fig. 11(a); if slit angles are nonzero to begin with, wedges account for the sum of the slit angle and the desired actuation angle. The octagonal panels are then taped to a flat surface with double-sided tape. We use a caliper to measure the overall cell extension (and, therefore, in-plane strains) and pop-up heights for various $\theta$ values. To account for specimen positioning and measurement inaccuracies, we perform each measurement three times, re-positioning the specimen each time. As a note, some plastic deformation and occasional tearing at the hinges is inevitable in these specimens.
D.2 Experiments on tube morphing mechanics
The tubes are subjected to tensile loading on an Instron 68FM-100 Universal Testing System (UTS) equipped with a 100 N load cell, at a loading rate of 1 mm/s. A picture of how a tube specimen is loaded onto the machine is shown in Fig. 11(b). Grippers are laser cut out of 6.4 mm-thick acrylic, and their shape is shown in Fig. 11(c). The top faces of the bottom anchor of the gripper are equipped with double sided tape. As shown in Fig. 11(d), the gripper is inserted into the end cap slot, then rotated and pulled once the bottom anchor is aligned with the engraved rectangle. Thanks to the tape, these grippers guarantee good adhesion to the tube end caps for tension experiments, at the low forces expected in our tests. For each tube, we carry out three measurements, from which we can compute an average and the standard deviation; the data is then reported as force-displacement curves in Fig. 6.
D.3 Experiments on tube-surface friction
To evaluate the frictional forces generated when tubes actuated at a desired displacement are dragged on a rough surface, we resort to the setup shown in Fig. 11(e). We manufacture an actuation bar, shown in Fig. 11(f), to keep the tube at a desired opening displacement, here chosen to be 2 cm as this is the maximum displacement we subject our tubes to in the morphing mechanics experiments. This rod is inserted into the tube as shown in Fig. 11(g), and it is designed so it can be gripped by the top clamp of our UTS; the other end of the rod is kept free. To perform frictional test, we take a strip of polyethylene foam and bond it to a 6.4 mm-thick acrylic plate. In turn, the plate is bonded to some fixtures to keep it vertical. The frictional surface is then placed at a distance $d$ from the axis of the tube, as shown in Fig. 11(e). This distance is chosen to be $d=3.2$ mm, which is slightly larger than the maximum half diameter of the octagonal cross section of our tubes; our choice is such that the acrylic end caps do not touch the frictional surface during our test, and the measured frictional force is only a product of the interactions between skin and surface. To perform a test, we displace the tube up by 3 cm and down by 3 cm, at a rate of 1 mm/s. When we test asymmetric skins, we expect the up ($F>0$) and down ($F<0$) tests to yield different results due to the asymmetry between pop-ups. Each test is repeated three times; each time, we reposition the frictional surface at the desired distance.
Appendix E Numerical modeling details
Here, we report some details on our numerical methods, and additional results that complement those in the main manuscript.
E.1 General model information